A tank can be filled in 4 days using pipe A alone, and in 3 days using both A and B. How many days will take to pipe B to fill the tank?
Answer by amoresroy(361)

A tank can be filled in 4 days using pipe A alone, and in 3 days using both A and B. How many days will take to pipe B to fill the tank?
Let b = number of days it will take pipe B to fill the tank
1/b= fraction of tank filled up per day using pipe b
1/4= fraction of tank filled up per day using pipe a
Since a tank can be filled in in 3 days using both A and B
(1/b +1/4)*3 = 1
3/b + 3/4 = 1
3/b = 1/4
b = 12
It will take 12 days for pipe B to fill the tank